OpenStudy (anonymous):

Two angles are Supplementary . One Angle is 6 more than four times the other. Find the measures of the two angles

OpenStudy (anonymous):

Suplementary means:\[\theta_1+\theta_2=180º\]One Angle is 6 more than four times the other means:\[\theta_1=6+4·\theta_2\]Replace this value of θ1in the first equation and you will have:\[6+4\theta_2+\theta_2=180 \rightarrow 5\theta_2=180-6=174 \rightarrow \theta_2=174/5=34.8\] Then θ1=180-34.8=145.2º Check that θ1=6+4θ2=6+4·34.8=145.2º